# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v2

@@ Abstract
Finding information in hour-long videos is a challenging task even for top-performing Vision Language Models (VLMs), as encoding visual content quickly exceeds available context windows. To tackle this challenge, we present FALCONEye [^1] Webpage: https://cplou99.github.io/FALCONEye/, a novel video agent based on a training-free, model-agnostic meta-architecture composed of a VLM and a Large Language Model (LLM). FALCONEye answers open-ended questions using an exploration-based search algorithm guided by calibrated confidence from the VLM’s answers. We also introduce the FALCON-Bench benchmark, extending Question Answering problem to Video Answer Search—requiring models to return both the answer and its supporting temporal window for open-ended questions in hour‑long videos. With just a 7B VLM and a lightweight LLM, FALCONEye outscores all open‑source 7B VLMs and comparable agents in FALCON-Bench. It further demonstrates its generalization capability in MLVU benchmark with shorter videos and different tasks, surpassing GPT‑4o on single‑detail tasks while slashing inference cost by roughly an order of magnitude.
@@ 1 Introduction
Vision Language Models (VLMs) have demonstrated remarkable performance on complex, fine-grained questions over images and short videos [llava-onevision, Qwen2.5-VL, openai2023gpt4v]. However, extending these capabilities to long-form videos remains an open challenge [chandrasegaran2024hourvideo, zhou2024mlvu, xiao2025videoqa]. Our work emphasis is on analyzing very long videos and single-detail questions whose answers are contained on a tiny set of frames, a needle in a haystack problem. Not only that, we go a step further by pushing the task boundaries—focusing on open-ended questions and requiring models to return both the answer and visual evidence of the answer in the form of a temporal window where the answer is found. This forces the agent to perform a deeper reasoning, limiting the number of hallucinations or random guesses deduced purely from the question or options without visual confirmation [rana2025mvu]. We refer to this extended setting as Video Answer Search (VAS) task.
Unlike general video understanding—where a global, embedded representation of the entire video may suffice for coarse-level tasks—VAS demands precise information retrieval within a large temporal space. Current VLMs typically address long video inputs by uniformly sampling frames (typically in the order of dozens or hundreds) until reaching the context window size limit [zhang2024llavanext-video, ye2024mplug]. While this strategy may support high-level summarization or scene classification, it often fails when the answer lies in a small, temporally-localized segment [xiao2025videoqa]. In VAS, locating a single relevant moment among thousands of frames requires not just compression, but active reasoning and targeted exploration—capabilities that current VLMs still lack.
Recent efforts have focused on fine-tuning VLMs for long-form video understanding [park2024too, wang2025videotree, shen2024longvu, song2024moviechat, longvila, Qwen2.5-VL, shu2025video, shu2025video]. These models aim to encode more frames by optimizing sampling and token compression strategies to fit within the LLM’s context window. While this enables a more complete global representation of the video, it comes at the cost of temporal resolution and precision—making them insufficient for tasks like VAS, where identifying a short, specific segment is crucial. To bridge this gap, several LLM-based agents such as such as VideoAgent [fan2024videoagent] or VideoTree [wang2025videotree] use meta-architectures combining multiple models to extract key frames—often via CLIP features—and answer multiple-choice questions.
@@
2) Novel video agent (FALCONEye): A system to emulate human-like behavior to answer single-detail questions in videos, integrating VLM and LLM capabilities in a training-free and model agnostic meta-architecture. It leverages our novel search algorithm, that reasons to iteratively focus on video clips most likely to contain the answer. FALCONEye and FALCON-Bench have been built under the lmms-eval framework [zhang2024lmmseval] to facilitate reproducibility and future evaluations. They are a significant advancement in VAS and, VQA in general, setting a new standard for precision-guided exploration in long-form video.
@@ 2 Related work
VLMs emerged to extend LLMs’ capabilities to vision, with the seminal work of LLaVA [liu2023llava], which integrated a pre-trained CLIP visual encoder (ViT-L/14 [radford2021learning]) to extract visual tokens, with an LLM (Vicuna [chiang2023vicuna]) for multimodal text and image understanding. This integration appears in other open-source architectures which improved the alignment and processing of visual features, such as BLIP-2 [li2023blip], LLaVA-1.5 [liu2023improvedllava], mPLUG-Owl2 [ye2024mplug], and LLaMa3.2-Multimodal [llama3_2].
More recently, Qwen2-VL [Qwen2-VL] introduced a VLM with a ViT capable of processing images at arbitrary resolutions, dynamically converting them into a variable number of visual tokens. Many companies have integrated increasingly powerful VLMs into their proprietary long-context chatbot systems, including OpenAI’s GPT-4V [openai2023gpt4v], Google DeepMind’s Gemini [google2023gemini] or Anthropic’s Claude [anthropic2024claude]. VLMs excel at image questions, but extending these capabilities to videos remains an open challenge due to the large number of tokens required to represent video content.
VLM strategies for video extend VLMs to process multiple frames as input [lin2023video, llava-next, li2024mvbench, llava-onevision, zhang2024llavanext-video] employing a uniform video sampling strategy to extract and process up to 64 frames, to keep the visual token amount within the LLM’s context window. Specific long-video solutions target efficient and sophisticated video representations. MovieChat [song2024moviechat] builds a global video representation combining short- and long-term token-based memories, Video-XL-Pro [liu2025video] and LongVU [shen2024longvu] present a spatiotemporal adaptive compression to retain the most relevant parts for question answering, and LongVILA [longvila] introduces Multi-Modal Sequence Parallelism to extend the number of video frames processed from 8 to 2048. Apollo [zohar2024apollo] conducts an extensive study on strategies to enhance long-video understanding. Differently, ReKV [di2025rekv] is a training-free mechanism that retrieves relevant information from a video key-value cache to enable efficient streaming VQA. Recently, Qwen2.5-VL [Qwen2.5-VL] extends its dynamic resolution capabilities to the temporal dimension, allowing processing up to 768 frames. Unlike these methods focused on answering by means of global complete video understanding, we focus on finding a short clip that contains the answer and returning the answer.
@@
ActivityNet-QA [yu2019activitynet] Human 5,800 58,000 180 O
MVBench [li2024mvbench] Auto 4,000 4,000 16 MC
EgoSchema [mangalam2023egoschema] Auto 5,031 5,031 180 MC
Event-Bench [du2024towards] Both 1,294 2,190 101 MC (21%)
LongVideoBench [wu2024longvideobench] Human 3,763 6,678 473 MC (30%)
LongVideoBench-L [wu2024longvideobench] Human 966 - 1408 MC (30%)
Video-MME [fu2024video] Human 900 2,700 1,017 MC (15%)
VIDEO MME-Long [fu2024video] Human 300 900 2466,7 MC (15%)
HourVideo [chandrasegaran2024hourvideo] Human 500 12,976 2,742 MC
LVBench [wang2024lvbench] Human 103 1,600 4,101 MC (14%)
NeXT-QA [xiao2021next] Human 5,440 52,044 44 MC/O
MovieChat-1K [song2024moviechat] Human 130 1,950 500 MC/O
MovieChat-1K $10^{\prime}$ [song2024moviechat] Human 19 - >600>600 MC/O
CG-Bench [chen2025cgbench] Human 1,219 12,129 1,624 MC/O (100%)
MLVU [zhou2024mlvu] Human 1,730 3,102 930 MC/O (64%)
MLVU $30^{\prime}$ [zhou2024mlvu] Human 125 - >1800>1800 MC/O (64%)
InfiniBench [ataallah2024infinibench] Auto 1,219 108,200 3,155 MC/O (14%)
QAEgo4D [Baermann_2022_qaego4d] Auto 1,325 14,513 495,1 MC/O (100%)
@@
VQA Benchmarks. Existing Video QA benchmarks evaluate VLMs capabilities for plenty of video comprehension tasks. Early benchmarks focus on short videos and questions to understand temporal actions [xiao2021next], human actions [yu2019activitynet], egocentric videos [mangalam2023egoschema] or movies [li2024mvbench]. Similar to ours, Event-Bench [du2024towards] focuses on single-detail events occurred within short temporal windows but uses one-minute-long videos. Recently, several VQA benchmarks [chandrasegaran2024hourvideo, wu2024longvideobench, zhou2024mlvu, fu2024video] extend the temporal duration of videos to around 10 minutes. Unlike these benchmarks, primarily focused on Multiple-Choice Questions (MCQs), MovieChat-1K [song2024moviechat] is centered on open-ended questions (OQs). LVBench [wang2024lvbench] and InfiniBench [ataallah2024infinibench] further push this challenge with longer videos, up to 1 hour. Some of these later benchmarks include a percentage of single-detailed questions, with the answer located within a short video time window, but only LongVideoBench [wu2024longvideobench] and InfiniBench evaluate a small fraction of the questions in open-ended mode. Recently, CG-Bench [chen2025cgbench] introduced a benchmark focused on single-detail questions and medium-duration videos (average of 26 minutes per video), being the first to provide ground-truth time intervals for answers and to evaluate state-of-the-art VLMs on the accuracy and temporal localization separately. QAEgo4D [Baermann_2022_qaego4d] also introduces a task close to VAS in egocentric videos based on the Ego4D NQL narrations and annotations [grauman2022ego4d]. However, it contains videos of average length 8 minutes and questions automatically generated from the narration, usually focused on the main actor or object being interacted. In contrast, our FALCON-Bench takes the challenge a step further: it moves to one-hour-long videos, focuses on open-ended questions, and evaluates the VAS task—requiring models to jointly return both the answer and its corresponding supporting temporal window. Table 1 highlights FALCON-Bench more challenging (longer and OE focused) benchmark for VAS task, including Human annotations.
@@
VQA Agent–style meta-architectures. Recent works have built agents for video QA combining an LLM with other architectures (e.g., VLMs, feature extractors). On one hand, there are fine-tuned solutions, e.g., TimeChat [ren2024timechat] fine-tunes a VLM with timestamp-aware encoder and sliding video Q-Former; or SeViLA [yu2023self] chains a fine-tuned BLIP-2 to localise frames before answering. On the other hand, there are very successful training-free pipelines. MoReVQA [min2024morevqa] and ProViQ [choudhury2024video] assemble multi-stage programs with ViperGPT [suris2023vipergpt] modules, while MVU [ranasinghe2024understanding] and VideoAgent2 [ranasinghe2024understanding] inject object-centric tracks into the LLM. Long Story Short [chung2023long] uses a VLM to caption clips, a LLM to select the most relevant ones for a given MCQ and visual-text cosine similarity from CLIP [radford2021clip] features to choose. the predicted option. Similarly, LifelongMemory [wang2024lifelongmemory] leverages an LLM to filter pre-extracted captions and answer MCQs from them. VideoAgent [wang2024videoagent] loops over (i) CLIP-based frame selection (ii) VLM captions of key frames, (iii) LLM answers a MCQ from captions with an uncalibrated confidence score until reaching a confident answer. VideoAgent[fan2024videoagent] couples a VLM captioner [zhao2023learning], a ViCLIP video encoder [wang2023internvid], a GPT‑based LLM, an object detector[zhao2024detrs] and an object tracker [zhang2022bytetrack] to tackle VQA. VideoTree [wang2025videotree] builds a hierarchical tree of key frames (pre-extracted with EVA CLIP-8B [sun2024eva]), captions each node with a VLM and queries the LLM for the final answer. The only agent that only requires a VLM and an LLM is the Socratic baseline [zeng2022socratic, zhang2023simple], where an LLM reasons only over VLM-generated captions. In this line, FALCONEye (i) is a training-free and model-agnostic meta-architecture that only requires a VLM and an LLM; (ii) is the first video agent to bring focus on strategies particularly designed for the more challenging open-ended VQA; (iii) its exploration algorithm is guided by a calibrated confidence estimate derived from the VLM’s open answer, not by an ad-hoc integer from the LLM; (iv) is powered by an adaptive search at clip-level: the agent varies clip length, frame count and resolution on-the-fly instead of relying on key-frame extraction ; (v) unlike existing agents, relying on dense captioning containing the answer, it reasons from very light captions to generate the answer directly from the target visual content.
@@ 3.2 Exploration Algorithm
Evaluation-. In this stage, each candidate clip $v^{*}_{i}$ is evaluated by the VLM. First, a uniform sampling of frames is performed within the clip. The number and resolution of sampled frames vary according to the duration of $v^{*}_{i}$, creating a zoom-in effect. For first-level clips, a larger number of frames with lower resolution are sampled, whereas for shorter clips, fewer frames with higher resolution are selected. For fixed input resolution VLMs, we use tiling to increase the resolution (similar to what GPT4o [openai2023gpt4v] does).
The VLM takes as input the question $Q$ along with the sampled frames from $v^{*}_{i}$ and generates a potential answer $\mathcal{A}^{*}_{i}$ and a confidence score $p(\mathcal{A}^{*}_{i})$. The answer is derived from a set of generated tokens $\mathcal{T}\in D$, such that $\mathcal{A}^{*}_{i}=Tokenizer(\mathcal{T})$, where $D=\{t_{1},\ldots,t_{d}\}$ is the token dictionary used by the LLM. During the VLM generation process, we store the logits $\{l_{1}^{k},\ldots,l_{d}^{k}\}$ for each generated token $t_{k}\in\mathcal{T}$ and compute $p(\mathcal{A}^{*}_{i})$ as the geometric average of the answer token probabilities [liu2023litcab]:
@@
where $s$ is the number of generated tokens in $\mathcal{T}$ (see Fig. 9 in supp. material). In MCQ scenario, there is no need for an average as $s=1$ (a single token for the option letter: A, B, C, D, …) and the confidence is just the softmax of the logits for the option letter token [chung2023long]. This step is supported by our VLM calibration analysis presented later in Section 6.

### v3

@@ Abstract
Finding information in hour-long videos is a challenging task even for top-performing Vision Language Models (VLMs), as encoding visual content quickly exceeds available context windows. To tackle this challenge, we present FALCONEye [^1] https://cplou99.github.io/FALCONEye/, a novel video agent based on a training-free, model-agnostic meta-architecture composed of a VLM and a Large Language Model (LLM). FALCONEye answers open-ended questions using an exploration-based search algorithm guided by calibrated confidence from the VLM’s answers. We also introduce the FALCON-Bench benchmark, extending Question Answering problem to Video Answer Search—requiring models to return both the answer and its supporting temporal window for open-ended questions in hour‑long videos. With just a 7B VLM and a lightweight LLM, FALCONEye outscores all open‑source 7B VLMs and comparable agents in FALCON-Bench. It further demonstrates its generalization capability in MLVU benchmark with shorter videos and different tasks, surpassing GPT‑4o on single‑detail tasks while slashing inference cost by roughly an order of magnitude.
@@ 1 Introduction
Vision Language Models (VLMs) have demonstrated remarkable performance on complex, fine-grained questions over images and short videos [22, 4, 36]. However, extending these capabilities to long-form videos remains an open challenge [6, 69, 56]. Our work emphasis is on analyzing very long videos and single-detail questions whose answers are contained on a tiny set of frames, a needle in a haystack problem. Not only that, we go a step further by pushing the task boundaries—focusing on open-ended questions and requiring models to return both the answer and visual evidence of the answer in the form of a temporal window where the answer is found. This forces the agent to perform a deeper reasoning, limiting the number of hallucinations or random guesses deduced purely from the question or options without visual confirmation [41]. We refer to this extended setting as Video Answer Search (VAS) task.
Unlike general video understanding—where a global, embedded representation of the entire video may suffice for coarse-level tasks—VAS demands precise information retrieval within a large temporal space. Current VLMs typically address long video inputs by uniformly sampling frames (typically in the order of dozens or hundreds) until reaching the context window size limit [65, 58]. While this strategy may support high-level summarization or scene classification, it often fails when the answer lies in a small, temporally-localized segment [56]. In VAS, locating a single relevant moment among thousands of frames requires not just compression, but active reasoning and targeted exploration—capabilities that current VLMs still lack.
Recent efforts have focused on fine-tuning VLMs for long-form video understanding [37, 54, 43, 45, 8, 4, 44, 44]. These models aim to encode more frames by optimizing sampling and token compression strategies to fit within the LLM’s context window. While this enables a more complete global representation of the video, it comes at the cost of temporal resolution and precision—making them insufficient for tasks like VAS, where identifying a short, specific segment is crucial. To bridge this gap, several LLM-based agents such as such as VideoAgent [18] or VideoTree [54] use meta-architectures combining multiple models to extract key frames—often via CLIP features—and answer multiple-choice questions.
@@
2) Novel video agent (FALCONEye): A system to emulate human-like behavior to answer single-detail questions in videos, integrating VLM and LLM capabilities in a training-free and model agnostic meta-architecture. It leverages our novel search algorithm, that reasons to iteratively focus on video clips most likely to contain the answer. FALCONEye and FALCON-Bench have been built under the lmms-eval framework [63] to facilitate reproducibility and future evaluations. They are a significant advancement in VAS and, VQA in general, setting a new standard for precision-guided exploration in long-form video.
@@ 2 Related work
VLMs emerged to extend LLMs’ capabilities to vision, with the seminal work of LLaVA [28], which integrated a pre-trained CLIP visual encoder (ViT-L/14 [38]) to extract visual tokens, with an LLM (Vicuna [9]) for multimodal text and image understanding. This integration appears in other open-source architectures which improved the alignment and processing of visual features, such as BLIP-2 [24], LLaVA-1.5 [27], mPLUG-Owl2 [58], and LLaMa3.2-Multimodal [1].
More recently, Qwen2-VL [49] introduced a VLM with a ViT capable of processing images at arbitrary resolutions, dynamically converting them into a variable number of visual tokens. Many companies have integrated increasingly powerful VLMs into their proprietary long-context chatbot systems, including OpenAI’s GPT-4V [36], Google DeepMind’s Gemini [12] or Anthropic’s Claude [2]. VLMs excel at image questions, but extending these capabilities to videos remains an open challenge due to the large number of tokens required to represent video content.
VLM strategies for video extend VLMs to process multiple frames as input [26, 23, 25, 22, 65] employing a uniform video sampling strategy to extract and process up to 64 frames, to keep the visual token amount within the LLM’s context window. Specific long-video solutions target efficient and sophisticated video representations. MovieChat [45] builds a global video representation combining short- and long-term token-based memories, Video-XL-Pro [29] and LongVU [43] present a spatiotemporal adaptive compression to retain the most relevant parts for question answering, and LongVILA [8] introduces Multi-Modal Sequence Parallelism to extend the number of video frames processed from 8 to 2048. Apollo [70] conducts an extensive study on strategies to enhance long-video understanding. Differently, ReKV [14] is a training-free mechanism that retrieves relevant information from a video key-value cache to enable efficient streaming VQA. Recently, Qwen2.5-VL [4] extends its dynamic resolution capabilities to the temporal dimension, allowing processing up to 768 frames. Unlike these methods focused on answering by means of global complete video understanding, we focus on finding a short clip that contains the answer and returning the answer.
@@
ActivityNet-QA [60] Human 5,800 58,000 180 O
MVBench [25] Auto 4,000 4,000 16 MC
EgoSchema [32] Auto 5,031 5,031 180 MC
Event-Bench [17] Both 1,294 2,190 101 MC (21%)
LongVideoBench [55] Human 3,763 6,678 473 MC (30%)
LongVideoBench-L [55] Human 966 - 1408 MC (30%)
Video-MME [19] Human 900 2,700 1,017 MC (15%)
VIDEO MME-Long [19] Human 300 900 2466,7 MC (15%)
HourVideo [6] Human 500 12,976 2,742 MC
LVBench [50] Human 103 1,600 4,101 MC (14%)
NeXT-QA [57] Human 5,440 52,044 44 MC/O
MovieChat-1K [45] Human 130 1,950 500 MC/O
MovieChat-1K $10^{\prime}$ [45] Human 19 - >600>600 MC/O
CG-Bench [7] Human 1,219 12,129 1,624 MC/O (100%)
MLVU [69] Human 1,730 3,102 930 MC/O (64%)
MLVU $30^{\prime}$ [69] Human 125 - >1800>1800 MC/O (64%)
InfiniBench [3] Auto 1,219 108,200 3,155 MC/O (14%)
QAEgo4D [5] Auto 1,325 14,513 495,1 MC/O (100%)
@@
VQA Benchmarks. Existing Video QA benchmarks evaluate VLMs capabilities for plenty of video comprehension tasks. Early benchmarks focus on short videos and questions to understand temporal actions [57], human actions [60], egocentric videos [32] or movies [25]. Similar to ours, Event-Bench [17] focuses on single-detail events occurred within short temporal windows but uses one-minute-long videos. Recently, several VQA benchmarks [6, 55, 69, 19] extend the temporal duration of videos to around 10 minutes. Unlike these benchmarks, primarily focused on Multiple-Choice Questions (MCQs), MovieChat-1K [45] is centered on open-ended questions (OQs). LVBench [50] and InfiniBench [3] further push this challenge with longer videos, up to 1 hour. Some of these later benchmarks include a percentage of single-detailed questions, with the answer located within a short video time window, but only LongVideoBench [55] and InfiniBench evaluate a small fraction of the questions in open-ended mode. Recently, CG-Bench [7] introduced a benchmark focused on single-detail questions and medium-duration videos (average of 26 minutes per video), being the first to provide ground-truth time intervals for answers and to evaluate state-of-the-art VLMs on the accuracy and temporal localization separately. QAEgo4D [5] also introduces a task close to VAS in egocentric videos based on the Ego4D NQL narrations and annotations [21]. However, it contains videos of average length 8 minutes and questions automatically generated from the narration, usually focused on the main actor or object being interacted. In contrast, our FALCON-Bench takes the challenge a step further: it moves to one-hour-long videos, focuses on open-ended questions, and evaluates the VAS task—requiring models to jointly return both the answer and its corresponding supporting temporal window. Table 1 highlights FALCON-Bench more challenging (longer and OE focused) benchmark for VAS task, including Human annotations.
@@
VQA Agent–style meta-architectures. Recent works have built agents for video QA combining an LLM with other architectures (e.g., VLMs, feature extractors). On one hand, there are fine-tuned solutions, e.g., TimeChat [42] fine-tunes a VLM with timestamp-aware encoder and sliding video Q-Former; or SeViLA [59] chains a fine-tuned BLIP-2 to localise frames before answering. On the other hand, there are very successful training-free pipelines. MoReVQA [33] and ProViQ [10] assemble multi-stage programs with ViperGPT [47] modules, while MVU [40] and VideoAgent2 [40] inject object-centric tracks into the LLM. Long Story Short [11] uses a VLM to caption clips, a LLM to select the most relevant ones for a given MCQ and visual-text cosine similarity from CLIP [39] features to choose. the predicted option. Similarly, LifelongMemory [53] leverages an LLM to filter pre-extracted captions and answer MCQs from them. VideoAgent [51] loops over (i) CLIP-based frame selection (ii) VLM captions of key frames, (iii) LLM answers a MCQ from captions with an uncalibrated confidence score until reaching a confident answer. VideoAgent[18] couples a VLM captioner [67], a ViCLIP video encoder [52], a GPT‑based LLM, an object detector[66] and an object tracker [64] to tackle VQA. VideoTree [54] builds a hierarchical tree of key frames (pre-extracted with EVA CLIP-8B [46]), captions each node with a VLM and queries the LLM for the final answer. The only agent that only requires a VLM and an LLM is the Socratic baseline [61, 62], where an LLM reasons only over VLM-generated captions. In this line, FALCONEye (i) is a training-free and model-agnostic meta-architecture that only requires a VLM and an LLM; (ii) is the first video agent to bring focus on strategies particularly designed for the more challenging open-ended VQA; (iii) its exploration algorithm is guided by a calibrated confidence estimate derived from the VLM’s open answer, not by an ad-hoc integer from the LLM; (iv) is powered by an adaptive search at clip-level: the agent varies clip length, frame count and resolution on-the-fly instead of relying on key-frame extraction ; (v) unlike existing agents, relying on dense captioning containing the answer, it reasons from very light captions to generate the answer directly from the target visual content.
@@ 3.2 Exploration Algorithm
Evaluation-. In this stage, each candidate clip $v^{*}_{i}$ is evaluated by the VLM. First, a uniform sampling of frames is performed within the clip. The number and resolution of sampled frames vary according to the duration of $v^{*}_{i}$, creating a zoom-in effect. For first-level clips, a larger number of frames with lower resolution are sampled, whereas for shorter clips, fewer frames with higher resolution are selected. For fixed input resolution VLMs, we use tiling to increase the resolution (similar to what GPT4o [36] does).
The VLM takes as input the question $Q$ along with the sampled frames from $v^{*}_{i}$ and generates a potential answer $\mathcal{A}^{*}_{i}$ and a confidence score $p(\mathcal{A}^{*}_{i})$. The answer is derived from a set of generated tokens $\mathcal{T}\in D$, such that $\mathcal{A}^{*}_{i}=Tokenizer(\mathcal{T})$, where $D=\{t_{1},\ldots,t_{d}\}$ is the token dictionary used by the LLM. During the VLM generation process, we store the logits $\{l_{1}^{k},\ldots,l_{d}^{k}\}$ for each generated token $t_{k}\in\mathcal{T}$ and compute $p(\mathcal{A}^{*}_{i})$ as the geometric average of the answer token probabilities [30]:
@@
where $s$ is the number of generated tokens in $\mathcal{T}$ (see Fig. 9 in supp. material). In MCQ scenario, there is no need for an average as $s=1$ (a single token for the option letter: A, B, C, D, …) and the confidence is just the softmax of the logits for the option letter token [11]. This step is supported by our VLM calibration analysis presented later in Section 6.